# Think-to-Talk or Talk-to-Think? When LLMs Come Up with an Answer in Multi-Hop Arithmetic Reasoning

### v3

@@ 4.2 Evaluation metrics
$t^{*}(v_{i})$ $=\min\{t\mid\max_{l}\mathrm{acc}(t,l,v_{i})>\tau\}\;\;\mathrm{,}$ (2)
where $\mathrm{acc}(t,l,v_{i})\in[0,1]$ indicates the accuracy of a probing classifier $f_{t,l,v_{i}}$. The $t^{*}(v_{i})\in\mathbb{Z}$ indicates when was the first time the probing classifier achieved a reasonable accuracy above $\tau$ (= 0.90 in our study [^4] See Appendix A.2 for results at different thresholds.) for the variable $v_{i}$. As a more coarse but comprehensive value, we also report $t_{\mathrm{eq}}^{*}(v_{i})$ indicating which equation $t_{\mathrm{eq}}$ the $t^{*}(v_{i})$ falls into. Given that the $t$ (and $t_{\mathrm{eq}}$) is relative to the CoT-beginning position, if $t_{\mathrm{(eq)}}^{*}(v_{i})$ is negative, the value $\{v_{i}\}$ is computed before CoT begins. This is the case for $t^{*}(i.e.,\mathrm{\texttt{B}})$ in Figure 2, based on the spike around $t_{\mathrm{eq}}$=-2 in the upper line graph.
We also report two types of accuracy:
@@
If $\mathrm{Acc}_{\mathrm{\prec CoT}}(v_{i})$ is sufficiently high, $\{v_{i}\}$ is resolved internally before CoT begins (Think-to-Talk mode). Conversely, if $\mathrm{Acc}_{\mathrm{\prec CoT}}(v_{i})$ is low and $\mathrm{Acc}_{\mathrm{\succ CoT}}(v_{i})$ is high, the answer is derived while performing CoT reasoning (Talk-to-Think mode).
4.3 Experimental results
Across task complexity levels.
We first analyze Qwen2.5-7B Qwen Team (2024) across the five task levels. Table 2 shows $t_{\mathrm{eq}}^{*}$ for each variable as well as the lower bounds of $t^{\dagger}_{\mathrm{eq}}$, which can be computed with a greedy resolver of equations. In most cases, regardless of #Steps or task level, $t^{*}_{\mathrm{eq}}>0$. The exceptions are $v_{2}$ in level 1 and $v_{3}$ in level 4. $v_{2}$ in level 1 has #Step =0=0 and is a variable whose value requires no computation to derive. $v_{3}$ in level 4 is a distractor and is not required to derive the final answer. Therefore, we find that the variables that are required to derive the final answer and require computation are all solved after CoT begins. In summary, we find that the model solved all subproblems necessary to derive the final answer during CoT, and that the Talk-to-Think mode is dominant.
@@ 4.3 Experimental results
Level variable #Step $t_{\mathrm{eq}}^{*}$ $t_{\mathrm{eq}}^{\dagger}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
1 $v_{1}$ ($\mathrm{A}$) 1 4 -2 0.36 1
$v_{2}$ ($\mathrm{B}$) 0 -2 -2 1 1
2 $v_{1}$ ($\mathrm{A}$) 1 2 -3 0.49 1
$v_{2}$ ($\mathrm{B}$) 2 5 -2 0.21 1
3 $v_{1}$ ($\mathrm{A}$) 2 5 -2 0.18 1
$v_{2}$ ($\mathrm{B}$) 1 2 -2 0.50 1
4 $v_{1}$ ($\mathrm{A}$) 2 5 -3 0.17 1
$v_{2}$ ($\mathrm{B}$) 1 2 -3 0.48 1
$v_{3}$ ($\mathrm{C}$) 1 N/A -2 0.44 0.24
5 $v_{1}$ ($\mathrm{A}$) 3 9 -2 0.18 1
$v_{2}$ ($\mathrm{B}$) 2 6 -2 0.23 1
$v_{3}$ ($\mathrm{C}$) 1 3 -2 0.51 1
Table 2: The results of Qwen2.5-7B on the five levels. The $t^{*}_{\mathrm{eq}}$ is the time when the model comes up with the correct answer (see §˜4.2). The $t_{\mathrm{eq}}^{\dagger}$ column indicates the lower bound of $t^{*}_{\mathrm{eq}}$ score. The $\mathrm{\prec CoT}$ and $\mathrm{\succ CoT}$ scores correspond to the accuracies introduced in §˜4.2. N/A indicates that the threshold $\tau$ was not exceeded at any position $t$.
@@
We further analyzed ten models listed in Table 3 on the Level 3 task. Same results as Qwen2.5-7B are generally obtained across various models, enhancing the generality of our obtained findings. These results are consistently observed across other tasks and irrespective of the threshold $\tau$. For more details, see §˜A.2.
@@
Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
@@
Table 3: Results for various models on the task Level 3. The $t^{*}$ column shows the token-wise time (described in §˜4.2), and the other columns are the same as Table 2. The $t^{*}$ and $t^{*}_{\mathrm{eq}}$ scores that are the same as their lower bounds are bolded.
@@
Figure 4: Success rates for each grid when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}}$) is the target token.
Figure 5: The upper part is the accuracy of probs, as shown in Figure 2. The lower part is the result of max pooling the Success rates from Figure 4 in the layer direction.
@@ 4.4 Analysis
In task Level 4 (see Table 1), $v_{3}$ (C) is a distractor, that is, $\{v_{3}\}$ is not necessary to derive the final answer. The models can infer this fact from the in context examples. According to Table 2, the $\mathrm{Acc}(v_{3})$ in Level 4 was at most 44%, a relatively low accuracy. From this result, we can see that, unlike the variables required to derive the final answer, $v_{3}$ is not encoded in a simple form that can be extracted by a linear transformation alone. This suggests the possibility that the model employs an efficient internal mechanism that does not derive variables unnecessary for obtaining the final answer. It is also consistent with the finding that a Talk-to-Think mode, in which computation is performed during CoT, is dominant.
@@ 5 Causal interventions
Motivation.
From our probing experiments in §˜4, we found that the sub-problems needed to produce the final answer are resolved after CoT-style generation begins. However, how this information propagates through the model remains unclear. To shed light on this flow of information, we intervene in the hidden states of the model during reasoning and analyze how these interventions affect the model’s outputs.
@@ 5.1 Settings
We employ activation patching Vig et al. (2020); Meng et al. (2022); Zhang and Nanda (2024), which is a widely adopted technique for causal intervention analysis. To inspect the causal relationship between specific hidden states $\bm{h}_{t,l}$ and a final answer $\hat{y}$, we compare two generation scenarios: (i) the ordinary inference and (ii) the intervened inference. In the latter scenario, we replace the specific hidden states $\bm{h}_{t,l}$ with other variants $\tilde{\bm{h}}_{t,l}$ obtained from the same model but with a different input $\tilde{x}$ (Clean run in Figure 3).
The input $x$ and $\tilde{x}$ have different correct answer $y$ and $\tilde{y}$ as well as different chains $z$ and $\tilde{z}$, respectively. For example, for the triple (x=x=``A=1+B,B=2+4;A=?,” z=z=``A=1+B,B=2+4,B=6,A=1+B,A=1+6,A=7,” y=y=7), one may use ($\tilde{x}$=``A=2+B,B=1+3;A=?,” $\tilde{z}$=``A=2+B,B=1+3,B=4,A=2+B,A=2+4,A=6,” $\tilde{y}$=6). If the model’s output turns from $y$ into $\tilde{y}$ or $z_{t}$ into $\tilde{z_{t}}$ due to the intervention to $\bm{h}_{t,l}$ with $\tilde{\bm{h}}_{t,l}$, we can confirm the causal relationship between $\bm{h}_{t,l}$ and the original answer $y$. We denote the model’s final output without intervention as $\hat{\tilde{y}}$ and that with intervention as $\hat{y}$, respectively ($\tilde{y}$ and $y$ denote respective gold answers). In the same way, we denote the generated reasoning chain without intervention as $\hat{\tilde{z_{t}}}$ and that with intervention as $\hat{z_{t}}$, respectively.
@@
We report Success Rate as a metric for this experiment. The Success rate indicates how frequently (%) the intervened output $\hat{\tilde{y}}$ aligns with the correct answer $\tilde{y}$. For reasoning chains, we report the Success rate for $\hat{\tilde{z_{t}}}$ as well.
@@
We specifically focus on Level 3 tasks and Qwen2.5-7B. Inspired by sliding window patching Zhang and Nanda (2024), we partition the hidden states into coarse grids, corresponding to each equation and every four layers, and perform activation patching on each grid separately (illustrated in Figure 5). [^5] All the hidden states in each grid are intervened at once. For every grid, we compute the Success rate by applying activation patching. We also examine multiple target tokens, specifically, at (i) the end of the equation 2 ($z_{17}$ in ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}\textbf{5}}}_{\hskip 1.42271pt\mathbf{2}}$ in Figure 2), (ii) the end of the equation 4 ($z_{32}$ in ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}\textbf{5}}}_{\hskip 1.42271pt\mathbf{4}}$ in Figure 2), and (iii) the final answer ($y$). When we apply activation patching, we generate the only target token with greedy decoding while forced-decoding the context. Note that the above equations are examples. We create a test set of 2,000 instances for evaluations.
@@
Figure 4 shows the Success rate for each grid when the final answer $y$ is the target token. We also show the results for the target tokens $z_{17}$ and $z_{32}$ in Appendix C. Figure 5 summarizes the max success rate among layers for each target token, and probing accuracy (same as the line graph in Figure 2).

### v4

@@ 3.2 Evaluation metrics
$t^{*}(v_{i})$ $=\min\{t\mid\max_{l}\mathrm{acc}(t,l,v_{i})>\tau\}\text{,}$ (2)
where $\mathrm{acc}(t,l,v_{i})\in[0,1]$ denotes the probing accuracy of $f_{t,l,v_{i}}$. The position $t^{*}(v_{i})\in\mathbb{Z}$ indicates when the first time the probing classifier achieved a reasonable accuracy above $\tau$. We set $\tau=0.9$ in our main experiment, and we show additional results with different thresholds in Appendix B.3. As a more coarse but comprehensive value, we also report $t_{\mathrm{eq}}^{*}(v_{i})$, indicating which equation $t_{\mathrm{eq}}$ in the input $x$ the position $t^{*}(v_{i})$ falls into. Given that the $t$ (and $t_{\mathrm{eq}}$) is relative to the CoT-beginning position, if $t_{\mathrm{(eq)}}^{*}(v_{i})$ is negative, the value $\mathdollar\{v_{i}\}$ is computed before CoT begins. In Figure 2, for example, both $t^{*}(\mathrm{\texttt{A}})$ and $t^{*}(\mathrm{\texttt{B}})$ are positive.
As complementary information of the position, we also report two types of accuracy:
@@
If $\mathrm{Acc}_{\mathrm{\prec CoT}}(v_{i})$ is sufficiently high, $\mathdollar\{v_{i}\}$ is resolved internally before CoT begins. Conversely, if $\mathrm{Acc}_{\mathrm{\prec CoT}}(v_{i})$ is low and $\mathrm{Acc}_{\mathrm{\succ CoT}}(v_{i})$ is high, the answer is derived while performing CoT reasoning.
3.3 Results
Across task levels.
We first analyze Qwen2.5-7B Qwen Team (2024) across the five task levels. Table 2 reports the position $t_{\mathrm{eq}}^{*}$ for each variable and its lower bound $t_{\mathrm{eq}}^{\mathrm{inf}}$ where the answer can be first identifiable. In most cases, (sub)answers are represented in the model’s hidden states corresponding to the Output part ($t^{*}_{\mathrm{eq}}>0$) in a linearly separable form, even when deriving relatively simple sub-answers, e.g., $\mathdollar\{\texttt{B}\}$ in Level 2. The exceptions are: (i) $v_{2}$ in Level 1, and (ii) $v_{3}$ in Level 4. Here, $v_{2}$ in Level 1 requires no computation (#Steps=0=0), and $v_{3}$ in Level 4 is a distractor that is not needed to derive the final answer. To sum up, all calculations (at least 1 step) required to obtain the final answer are resolved during CoT, suggesting that it’s unlikely to refer to a predetermined answer as model’s final answer.
@@ 3.3 Results
Level variable #Step $t_{\mathrm{eq}}^{*}$ $t_{\mathrm{eq}}^{\mathrm{inf}}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
1 $v_{1}$ ($\mathrm{A}$) 1 4 -2 35.8 100
$v_{2}$ ($\mathrm{B}$) 0 -2 -2 100 100
2 $v_{1}$ ($\mathrm{A}$) 1 2 -3 49.2 100
$v_{2}$ ($\mathrm{B}$) 2 5 -2 21.6 94.7
3 $v_{1}$ ($\mathrm{A}$) 2 5 -2 17.9 97.4
$v_{2}$ ($\mathrm{B}$) 1 2 -2 50.5 100
4 $v_{1}$ ($\mathrm{A}$) 2 5 -3 17.2 100
$v_{2}$ ($\mathrm{B}$) 1 2 -3 47.7 100
$v_{3}$ ($\mathrm{C}$) 1 N/A -2 43.7 23.7
5 $v_{1}$ ($\mathrm{A}$) 3 9 -2 18.1 100
$v_{2}$ ($\mathrm{B}$) 2 6 -2 22.6 100
$v_{3}$ ($\mathrm{C}$) 1 3 -2 50.6 100
Table 2: The results of Qwen2.5-7B on the five levels. The $t^{*}_{\mathrm{eq}}$ is the time when the model comes up with the correct answer (see §˜3.2). The $t_{\mathrm{eq}}^{\mathrm{inf}}$ column indicates the lower bound of $t^{*}_{\mathrm{eq}}$ score. The $\mathrm{\prec CoT}$ and $\mathrm{\succ CoT}$ scores correspond to the accuracies introduced in §˜3.2. N/A indicates that the threshold $\tau$ was not exceeded at any position $t$.
@@
We analyzed nine models (Table 3) on the Level 3 task. Similarly to Table 2, we observed $t^{*}_{\mathrm{eq}}>0$. This pattern holds for other tasks and for different thresholds $\tau$ (Tables 923 in §˜B.3). On the other hand, in the Input part—particularly the latter half of equation 2-2 ($\underline{\mathrm{B}=2+3,}_{\hskip 1.42271pt\mathbf{-2}}$)—the pre-CoT accuracy $\mathrm{Acc}_{\mathrm{\prec CoT}}(v_{i})$ increases modestly, peaking around $60\%$.
@@
Var. $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
@@
Table 3: Results for various models on the Level 3 task. The $t^{*}$ column shows the token-wise time (described in §˜3.2), and the other columns are the same as Table 2. The $t^{*}$ and $t^{*}_{\mathrm{eq}}$ scores that are the same as their lower bounds are bolded.
Figure 3: Error analysis of cases where Llama3.2-3B generated incorrect answers. The vertical axis represents the index of the transformer layer. The horizontal axis represents the tokens input to the model over time. The numbers highlighted in green represent the gold labels for the predictions, while those highlighted in red denote the values incorrectly generated by the model.
@@
Figure 5: Results of the causal intervention on Qwen2.5-7B. Each grid cell shows the success rate when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}})$ is the target token.
Figure 6: The upper part is the accuracy of probes, as shown in Figure 2. The lower part is the result of max pooling the Success rates from Figure 5 in the layer direction.
@@ 3.4 Analysis
In this section, we present the results of a prompt-sensitivity analysis and an error analysis. We also analyze the effects of factors such as reasoning-chain format (including implicit reasoning). For details, see appendix˜A.
Equation order.
The difference between these two tasks lies in the order of the equations. Level 3 tasks involve forward references, whereas Level 2 tasks do not. An autoregressive language model can attend only to past inputs. Therefore, the model may be able to reach the final answer more quickly on Level 2 tasks than on Level 3 tasks. In practice, Table 2 indicates that the probe predicts the answer during CoT at both levels. Therefore, the impact of equation order (dependency structure) on the model’s internal mechanism is thought to be limited.
@@
In the Level 4 task (see Table 1), the variable $v_{3}$ (C) is a distractor, that is, $\mathdollar\{v_{3}\}$ is not necessary to derive the final answer. The models can infer this fact from the few-shot examples. According to Table 2, the $\mathrm{Acc}(v_{3})$ in Level 4 was at most 44%, a relatively low accuracy. From this result, we can see that, unlike the variables required to derive the final answer, $v_{3}$ is not encoded in a simple form that can be extracted by a linear transformation alone. This suggests the possibility that the model employs an efficient internal mechanism that does not derive variables unnecessary for obtaining the final answer. It is also consistent with the finding that the CoT is faithful to the final answer.
Effect of depth of reasoning.
The difference between Level 3 and Level 5 is whether two or three calculation steps are required to reach the final answer. Qualitatively, the probe’s prediction accuracy for each variable increases a few tokens before the answer appears in the output text, indicating alignment between the internal state and the output. This finding supports the conclusion that the generated explanations are faithful to the internal state and that this conclusion generalizes across levels.
Variable When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Setting Variable #Step $t_{\mathrm{eq}}^{*}$ $t_{\mathrm{eq}}^{\dagger}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Same level $v_{1}$ 2 5 -3 17.2 100
$v_{2}$ 1 2 -3 47.7 100
$v_{3}$ 1 N/A -2 43.7 23.7
General $v_{1}$ 2 5 -3 18.1 94.7
$v_{2}$ 1 2 -3 49.4 100
$v_{3}$ 1 N/A -2 43.3 28.1
Table 4: Evaluation results of same level prompting and general prompting. Each column is the same as Table 2.
Prompt differences.
We investigated whether the format of few-shot prompting influences the model’s internal reasoning patterns. We analyzed the differences between two scenarios: 1. Same level prompting and 2. General prompting. In same level prompting, we provide three equations that match the level of the evaluation task, following the general setting (§˜2). In contrast, under general prompting the model receives 50 randomly generated demonstrations (three equations each). [^4] We increased the number of demonstrations in the general prompts because task performance was otherwise insufficient. Table 4 reports $t_{\mathrm{eq}}^{*}$ for both strategies for Level 4 task. We observe no difference in $t_{\mathrm{eq}}^{*}$, suggesting that the few-shot prompt format has minimal effect on the model’s internal reasoning patterns.
Behavior in incorrect instances.
Figure 3 shows the top-1 predictions of the probe for $\mathrm{B}$ in instances where Llama3.2-3B produced incorrect answers for level 3 (where B happened to be 4). When the model produced an incorrect answer, we often found that the correct answer had appeared at an earlier decoding step (see §˜B.2 for the probe’s predictions on additional samples). The underlying reason for these mistakes remains unclear, but through our probing framework, one can track how such an incorrect answer is propagated to the subsequent generation, which might be practically helpful in debugging the model internals.
@@ 4 Causal interventions
From our probing experiments in §˜3, we tentatively concluded that the final answer is determined after CoT-style generation begins. That is, CoT is faithful to the final answer. In this section, we corroborate this conclusion using causal intervention analysis and further clarify how input information flows internally while generating CoT reasoning and the final answer.
@@ 4.1 Settings
We employ activation patching Vig et al. (2020); Meng et al. (2022); Zhang and Nanda (2024), which is a widely adopted technique for causal intervention analysis. To inspect the causal relationship between specific hidden states $\bm{h}_{t,l}$ and a final answer $y$, we compare two generation scenarios: (i) the ordinary inference and (ii) the intervened inference. In the latter scenario, we replace the specific hidden states $\bm{h}_{t,l}$ with other variants $\tilde{\bm{h}}_{t,l}$ obtained from the same model but with a different input $\tilde{x}$ (Intervention in Figure 4).
The input $x$ and $\tilde{x}$ have different correct answer $y$ and $\tilde{y}$ as well as different chains $z$ and $\tilde{z}$, respectively. For example, for the triple ($\tilde{x}$=``A=1+B,B=2+4;A=?,” $\tilde{z}$=``A=1+B,B=2+4,B=6,A=1+B,A=1+6,A=7,” $\tilde{y}$=7), one may use (x=x=``A=2+B,B=1+3;A=?,” z=z=``A=2+B,B=1+3,B=4,A=2+B,A=2+4,A=6,” y=y=6). If the model’s output turns from $y$ into $\tilde{y}$ or $z_{t}$ into $\tilde{z_{t}}$ due to the intervention to $\bm{h}_{t,l}$ with $\tilde{\bm{h}}_{t,l}$, we can confirm the causal relationship between $\bm{h}_{t,l}$ and the original answer $y$. Henceforth, to precisely explain the setting, we denote the model’s final output with intervention as $\hat{y}^{\mathrm{patch}}$ and that without intervention as $\hat{y}$, respectively ($y$ and $\tilde{y}$ denote respective gold answers). In the same way, we denote the generated reasoning chain with intervention as $\hat{z_{t}}^{\mathrm{patch}}$ and that without intervention as $\hat{z_{t}}$, respectively.
@@
We report Success Rate as a metric for this experiment. The Success rate indicates how frequently (%) the intervened output $\hat{y}^{\mathrm{patch}}$ aligns with the correct answer $\tilde{y}$. For reasoning chains, we report the Success rate for $\hat{z_{t}}^{\mathrm{patch}}$ as well.
@@
We specifically focus on Level 3 tasks and Qwen2.5-7B. The experimental results for the other models are reported in §˜B.4. Inspired by sliding window patching Zhang and Nanda (2024), we partition the hidden states into coarse grids, corresponding to each equation and every four layers, and perform activation patching on each grid separately (illustrated in Figure 6). [^5] All the hidden states in each grid are intervened at once. For every grid, we compute the Success rate by applying activation patching. We also examine multiple target tokens, specifically, at (i) the end of the equation 2 ($z_{17}$ in ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}\textbf{5}}}_{\hskip 1.42271pt\mathbf{2}}$ in Figure 2), (ii) the end of the equation 4 ($z_{32}$ in ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}\textbf{5}}}_{\hskip 1.42271pt\mathbf{4}}$ in Figure 2), and (iii) the final answer ($y$). When we apply activation patching, we generate only the target token with greedy decoding while forced-decoding the context. Note that the above equations are examples. We create a test set of 2,000 instances for evaluation.
@@
CoT is faithful to the final answer.
Figure 5 presents the success rate at each grid position when the final answer $y$ (${\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{2}}$) is the target token. The results show a strong causal dependence of the final answer on the Output part, whereas the Input part has only a limited effect. We observe the same pattern for other target tokens (e.g., $z_{17}$ and $z_{32}$). This trend is broadly consistent across models (§˜B.4). Furthermore, we focus on the generation of ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}\textbf{5}}}_{\hskip 1.42271pt\mathbf{2}}$, which is a necessary subanswer to reach the final answer (see Figure 6). Intervention to the immediately preceding equation ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=2+3}}}_{\hskip 1.42271pt\mathbf{1}}$ affects the generation of ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}\textbf{5}}}_{\hskip 1.42271pt\mathbf{2}}$, but patching the equation ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=2+3}}}_{\hskip 1.42271pt\mathbf{-2}}$ in the problem statement $x$ (the region where a slight improvement in probing accuracy before the CoT was observed; t=11t=-11 to t=4t=-4) had almost no effect on the output of ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}\textbf{5}}}_{\hskip 1.42271pt\mathbf{2}}$. Similarly, when we focus on the case where the target token is ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}\textbf{5}}}_{\hskip 1.42271pt\mathbf{4}}$, patching the equation ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=2+3}}}_{\hskip 1.42271pt\mathbf{-2}}$ again resulted in almost no change in the output. Therefore, the influence of the information processed in the Input part on the sub-answers (${\color[rgb]{.5,.5,.5}\mathrm{B}=}\textbf{5}$) is limited. This tendency is especially strong in the Qwen family (except for Qwen2.5-Math-7B). By contrast, other model families show a weak causal relationship between ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}\textbf{5}}}_{\hskip 1.42271pt\mathbf{2}}$ and the Input part (§˜B.4). This aligns with the results observed in probing experiments §˜3.3.
Recency bias.
@@
5 Related Work
Post hoc nature of CoT.
Several prior studies investigate similar questions Afzal et al. (2025); Ye et al. (2025); Cox (2025); Lewis-Lim et al. (2025). For instance, Cox (2025) examined whether CoT explanations are post-hoc or not, specifically, confirmed that, in a binary classification task, the correct label can be predicted from the hidden state at the end of the input (prompt). In their setting, the CoT $z$ is (freely) produced from the intervened hidden state $x$, while we force-decode the CoT chain. In this sense, their scope is rather on the causality of $x\rightarrow z$, and thus, as stated in §˜2.3, the scope is slightly different. They also showed that, in some tasks (e.g., sports understanding), the final answer can be predicted with the model’s final state in the prompt, apparently contradicting our probing experiments, but their tasks would be answerable with factual or commonsense knowledge without CoT, and in their specific logical deduction task—where on-the-fly reasoning is required, similarly to our arithmetic tasks—the probe indeed struggles to predict the final answer, similarly to our results. Afzal et al. (2025) also suggests the model reaches the answer internally before CoT begins, but their experimental design is somewhat different from ours. Specifically, they try to predict whether the model will succeed/fail the task before CoT begins as a binary problem, rather than tracing how models reach the correct answer during CoT or identifying where the concrete answer comes up with during CoT reasoning. Ye et al. (2025) also find that the model, in advance, anticipates which components (variables) will be required to reach the final answer by tracking whether a variable has been computed, and how models retain computed values during reasoning. Although the experiments are similar, they do not investigate causal relationships and therefore do not fully answer the question of whether the reasoning chain is faithful.
Faithfulness of CoT.
Paul et al. (2024); Bentham et al. (2024) similarly evaluate the faithfulness of CoT to the final answer. They test the consistency between the final answer and the CoT by intervening on the CoT text. Specifically, they replace it with counterfactual content and then check whether the final answer changes. They report that CoT is not sufficiently faithful to the final answer. In addition, Turpin et al. (2023); Chen et al. (2025) define CoT as faithful when the following two conditions hold: (i) The CoT explicitly refers to the hint provided in the prompt. (ii) The model relies on the hint to reach the answer (that is, without the hint, the model would not have produced the hinted answer). Using these conditions, they evaluate faithfulness indirectly by testing whether the model’s internal computation depends on the hinted information.
By contrast, in our work, we trace the model’s internal reasoning more directly by probing the values of intermediate answers, and thus track the faithfulness of CoT to the final answer using a more direct method and a controlled testbed than these prior studies. Through token-level probing, we analyze when and at which token position the reasoning is actually carried out. In this respect, our work differs in that it also allows us to observe the faithfulness of the internal states to the CoT text along the temporal axis.
Interpreting multi-hop reasoning in language models.
Interpreting internal mechanisms of LLMs has been actively investigated Conneau et al. (2018); Tenney et al. (2019); Niven and Kao (2019); nostalgebraist (2020); Geva et al. (2023); Lieberum et al. (2024); Ghandeharioun et al. (2024); Ferrando and Voita (2024). They revealed, for example, specialized attention heads responsible for specific operations (Cabannes et al., 2024) or decision-making, copying, and induction (Dutta et al., 2024). With a more concrete example, Yang et al. (2024c) showed that, even during the first pass of the problem statements such as The mother of the singer of Thriller is ___, language models first resolve a bridge entity, Stevie Wonder in this case, then identify the final answer. This study is more focused on the difference between the first pass of the problem statements (before CoT generation) and their second pass involving explicit problem solving (while CoT generation).
Arithmetic representations in LLMs.
How models handle numerical information has also been closely studied. For instance, Heinzerling and Inui (2024) used partial least squares regression Wold et al. (2001) to demonstrate that numeric attributes, such as birth years and population numbers, are encoded as monotonic representations in the activation space of LLMs and can be manipulated with interventions. In turn, Stolfo et al. (2023) showed that, in autoregressive LLMs, the operations and numerical information necessary for solving quantitative reasoning are processed in the lower layers of the model, and these results are used by the attention layers to predict the final calculation outcomes. Zhu et al. (2025) studied the representation of numbers in language models’ hidden states during single-hop arithmetic tasks (e.g., What is the sum of 12 and 34?). Their analysis revealed that numerical information is encoded linearly within the hidden states and demonstrated that individual digits could be manipulated independently. In this study, we add to this literature by introducing incremental arithmetic problem solving, i.e., what numerical information is contained in the model’s hidden states at each time step of multi-hop arithmetic reasoning.
Model interpretability methods.
Linear probing Alain and Bengio (2017b) is one of the representative methods for analyzing the internal representations of neural models—a small model predicts a specific feature from them, thereby determining whether the input contains information about that feature. In this study, we use them to derive the models’ intermediate answers. The causality with the model’s output can be further verified by examining if a model’s predictions change when the hidden states are intervened Li et al. (2023); Wu et al. (2023). One representative intervention method is activation patching Vig et al. (2020); Meng et al. (2022); Zhang and Nanda (2024), where hidden states obtained from one model instance are transplanted onto another during inference to change its predictions. Such techniques can be applied as a way to control model behavior in practical scenarios such as mitigating inherent biases Zhao et al. (2019); Ganguli et al. (2023); Yang et al. (2024b). Here, we employ activation patching to validate the plausibility of the probing results.